OpenStudy (anonymous):

Which lines has a slope of -1/2? A.)x + 2y = 0 B.)x - 2y = 0 C.)-x + 2y = 0

OpenStudy (anonymous):

B

OpenStudy (raden):

if given a straight line ax+by=c then the slope always : m = -a/b

OpenStudy (raden):

B is incorrect

OpenStudy (anonymous):

yea i found my error

OpenStudy (whpalmer4):

Another form that is useful is slope-intercept form: \[y=mx+b\] where \(m\) is slope and \(b\) is y-intercept value. I find this easier to remember, but your mileage may vary.

OpenStudy (whpalmer4):

nothing worse than getting the problem wrong because you remembered the formula with an incorrect sign, or some terms reversed, or some other dumbass mistake like that :-)

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

[Find the slope of the line whose equation is 5y = x - 3.

OpenStudy (anonymous):

Well it's clearly A.

OpenStudy (whpalmer4):

Okay, we'll go through a few of them. \[-x+2y=0\]\[2y=x\]\[y=x/2=\frac{1}{2}x\]slope is 1/2 \[x+2y = 0\]\[2y=0-x\]\(y=-x/2 =-\frac{1}{2}x\) That means the slope is -1/2

OpenStudy (anonymous):

I understand the answer but I just dont get the y=mx+b i never did

OpenStudy (whpalmer4):

You don't get why it works, or how I got the equation into that form, or what?

OpenStudy (anonymous):

I dont get how to use it and why.

OpenStudy (anonymous):

AAAAAAAAAAAAAAAAAAAA I have done this recently

OpenStudy (anonymous):

You all are over complicating it

OpenStudy (whpalmer4):

by showing you how to work the problem, we are over-complicating it? okay, whatever...be sure to let us know what the easy way is :-)

OpenStudy (anonymous):

Yea but your not even figuring it out or it would have already been solved.